7.16 A golf ball of mass 0.045 kg is hit off the tee at a speed of 33 m/s . The golf club was in contact with the ball for 3.90×10−3 s Find the impulse imparted to the golf ball. Find the average force exerted on the ball by the golf club.
Answer #1

given that

mass of a ball m = 0.045 kg

speed u = 33 m/s

time t = 3.9*10^-3 s

from the impulse momentum theorem

change in momentum = impulse

impulse = m(u-0)

impulse = f avg*t = mu

f avg = 0.045*33/3.9*10^-3

f avg = 380.7 N